Find the equation of the circle that has its center at (1,2) and that passes through the origin.
Answer by Alan3354(69443)

Find the equation of the circle that has its center at (1,2) and that passes through the origin.
-------------------
The radius is the distance from the center to the origin.
r^2 = 1^2 + 2^2 = 5
-------------
(x-1)^2 + (y-2)^2 = 5